OpenStudy (anonymous):

200m race. A wins B by 10m or 2 sec. Speed of A and B

OpenStudy (ghazi):

lets say B took t seconds, then time taken by A is = t-2 then Speed of A= 200/ (t-2) and speed of B = 200/t

OpenStudy (anonymous):

i dont see any point in wining race by 10m lol both have to complete 200m